OpenStudy (anonymous):

How much would you need to deposit in an account now in order to have $20,000 in the account in 4 years? Assume the account earns 5% interest.

OpenStudy (muscrat123):

do u think it will b more or less than 20,000?

OpenStudy (david27):

what is b

OpenStudy (muscrat123):

b = be

OpenStudy (david27):

I think less

Vocaloid (vocaloid):

A = P(1+r)^t let A = 20,000, r = 0.05, and t = 4 solve for P

Vocaloid (vocaloid):

more specifically... (total money) = (money deposited)(1+interest/100)^time

OpenStudy (david27):

vocaloid

OpenStudy (wolf1728):

If you need more time, let's change vocaloid's equation a little bit beginning money = [total money] / [(1.05)^4] Can you solve that now?

OpenStudy (anonymous):

It's stil unclear

OpenStudy (wolf1728):

Do you know how to raise a number to a power? Part of that equation is asking you to raise 1.05 to the 4th power. Can you do that?

OpenStudy (anonymous):

Yea

OpenStudy (wolf1728):

What is 1.05 to the 4th power?

OpenStudy (anonymous):

1.215

OpenStudy (wolf1728):

Yes So according to the equation, you get 20,000 and divide it by 1.215

OpenStudy (wolf1728):

To be even more exact, 1.05^4 = 1.21550625 and so the starting money would be 16,454.05